OpenStudy (anonymous):

x and y are real x^2+2xy-y^2=6. find minimum value of (x^2+y^2)^2

OpenStudy (amistre64):

i wonder, the left looks like a complete square already, or almost x^2+2xy-y^2=6 x^2+2xy+(y^2-y^2)-y^2=6 (x+y)^2 -2y^2=6 (x+y)^2 = 6+2y^2

ganeshie8 (ganeshie8):